OpenStudy (anonymous):

In the process of developing a photographic film, silver bromide is dissolved by adding a solution of sodium thiosulphate: AgBr(s)+2NaS2O3=Na3Ag(S2O3)2(aq)+NaBr(aq) If you want to dissolve0.250g of AgBr (fw 187.8), how many milliliters of 0.0139 M Na2S2O3 should you add?

OpenStudy (anonymous):

I know the right answer is 163mL. I just can't figure out how to get that answer. I'm not even getting close to the right answer.

OpenStudy (anonymous):

Sorry that is 193mL NOT 163mL

OpenStudy (aaronq):

since everything stoichiometrically reacts in ratios of moles, convert all the units into moles

OpenStudy (aaronq):

for every mole of AgBr you need 2 of NaS... then once you know the amount of moles molarity=moles/L of solvent

OpenStudy (anonymous):

Yes but I'm not getting to 193mL. Im getting really small decimal amounts

OpenStudy (aaronq):

did you convert from L to millilitres?

OpenStudy (anonymous):

ok I got it thank you. Its 0.250g(1/187.8g)=1.00133molAgBr x (2/1)=0.00266molNa2S2O3 0.0138M=0.0138mol/L x(1L/1000mL)=0.0000138mol/mL 0.00266/0.0000138=193mL Na2S2O3